OpenStudy (anonymous):

solve the 2nd order non-homogeneous linear recurrence relation with constant coefficients: f(n+2)+f(n+1)-6f(n)=-8n+22 subject to the start-up conditions f(0)=3; f(1)=2

OpenStudy (anonymous):

we will let you solve this one

OpenStudy (anonymous):

i wish i knew how to solve it ..i wouldn't be asking for help if i knew

OpenStudy (anonymous):

look at previous one ;solved by others for you

OpenStudy (anonymous):

i know i am looking at it and trying to make sense of it

OpenStudy (anonymous):

hi mathmate

OpenStudy (anonymous):

ok so the first step is \[r ^{2}+r-6=8n+22 \]?

OpenStudy (anonymous):

r= -3,2 ?

OpenStudy (anonymous):

solve for homogenous first

OpenStudy (anonymous):

then plug in to solve for particular case

OpenStudy (anonymous):

hold on let me try to figure out

OpenStudy (mathmate):

I'm here, but not finished looking at the numbers.

OpenStudy (anonymous):

okay no problem :) ..im also trying to figure out how to do this by looking at the previous problems

OpenStudy (mathmate):

OK, I've got the solution. It takes a little while because we can convert the given inhomogeneous relation to a 4th order homogeneous, which we know how to solve easily.

OpenStudy (mathmate):

It's a little more complicated than that because of the right-hand side.

OpenStudy (mathmate):

My aim is to eliminate the rhs by elimination, and it turns out that it takes two steps.

OpenStudy (mathmate):

right-hand-side.

OpenStudy (mathmate):

What we do, is to increase n by 1, and write the relation as f(n+3)+f(n+2)-6f(n+1)=-8(n+1)+22

OpenStudy (mathmate):

Subtract the original relation from it to get: f(n+3)-7f(n+1)+6f(n)=-8

OpenStudy (mathmate):

We still have 8 on the rhs, so we do the same thing again to get f(n+4)-f(n+3)-7f(n+2)+13f(n+1)-6f(n)=0 You're OK with that so far?

OpenStudy (anonymous):

hold on im still reading the second part u wrote

OpenStudy (anonymous):

where did u get the 7 from in 7f(n+1)?

OpenStudy (anonymous):

and what happened to the 22?

OpenStudy (mathmate):

When you subtract the first equation from the second, the 22 disappears, and the -7f(n+1) comes from -6f(n+1)-f(n+1).

OpenStudy (mathmate):

If it helps, give me 4-5 minutes, I will write it up properly so it will be easier to follow. Is that OK?

OpenStudy (anonymous):

oh sure it's no problem at all thank yyou

OpenStudy (mathmate):

Will be right back.

OpenStudy (anonymous):

please take your time :)

OpenStudy (mathmate):

Question: to solve f(n+2)+f(n+1)-6f(n)=-8n+22 ...(1) given f(0)=3, f(1)=2. Since the equation is non-homogeneous, we will try to eliminate the right-hand-side by elimination, so that the characteristic equation method will work. The elimination is done by increasing n, and subtracting. Substitute n+1 for n in (1) gives f(n+3) + f(n+2) - 6f(n+1) = -8(n+1) + 22 ....(2) Subtract (1): f(n+2) + f(n+1) - 6f(n) = -8(n) + 22 ....(1) to get f(n+3) - 7f(n+1) + 6f(n) = 8 .........(3) Now, repeat the same process to eliminate 8: f(n+4) - 7f(n+2) + 6f(n+1) = 8 ...(3A) subtract: f(n+3) - 7f(n+1) + 6f(n) = 8 ....(3) to get f(n+4)-f(n+3)-7f(n+2)+13f(n+1)-6f(n) = 0 ..,.(4) Equation (4) is a homogeneous equation which can be solved by the characteristic equation method. Solving 4 gives the roots as x=1,1,2,3 (double root at 1). So the general solution is: f(n)=(c1 + c2 n) + c3.2^n + c4.(-3)^n To solve for c1-c4, we need 4 values, of which we were given two: f(0)=3 f(1)=2 so calculate from equation (1) f(2)=22 f(3)=-12 f(4)=134 (for checking later on). Set up a linear system by substituting f(0)=3=c1+c2(0)+c3+c4 f(1)=2=c1+c2+2C3-3C4 f(2)=22=c1+2c2+4c3+9c4 f(3)=-12=c1+3c2+8c3-27c4 Solve the system for c1-c4 as [0,2,9/5,6/5] which means that the solution is: f(n)=2n+(9/5)2^n+(6/5)(-3)^n Check that f(n) works for at least n=0-4.

OpenStudy (anonymous):